# Game-Theoretic Defenses for Robust Conformal Prediction Against Adversarial Attacks in Medical Imaging

### v1

@@ 5.2.3. Results for RQ3
In the third experiment, an unknown and potentially adversarial attack is applied to the test data. The prediction set size corresponds to the attacker’s payoff and the defender’s cost, framing the interaction as a zero-sum game. To develop an optimal solution for the defender in a strategic manner, we preserve an additional evaluation set, which is exchangeable with both calibration and test sets. We use the prediction set size evaluated on this evaluation set as the estimated payoff matrix and derive the Nash equilibrium to determine the defender’s optimal strategy. This strategy is then applied during the test phase.
@@
The payoff matrix in this scenario has a shape of $6\times 3$, accounting for the five defensive strategies (including $f_{\text{max}}$ and $f_{\text{min}}$) and the three attack types. Examples of pairs of estimated payoff matrices and their corresponding true payoff matrices are displayed in Figures 2, 3, and 4 for each of the three datasets.
We evaluate the performance of different defensive strategies, including the optimal strategy derived from the Nash equilibrium on the evaluation set, as well as simple strategies and a uniform strategy that assigns equal probability to each defensive model. The prediction set sizes on the test set using these strategies are reported in Figures 5(a), 5(b), and 5(c). The results indicate that the optimal strategy consistently outperforms other strategies, often degenerating to a singular defensive strategy rather than a mixed strategy, thereby achieving the best performance across different datasets.
Figure 2. RQ3: Estimated and True Payoff Matrices for PathMNIST
Figure 3. RQ3: Estimated and True Payoff Matrices for OrganAMNIST
Figure 4. RQ3: Estimated and True Payoff Matrices for TissueMNIST
(a) PathMNIST
(b) OrganAMNIST
(c) TissueMNIST
Figure 5. RQ3: Comparison of Strategies Across Different Datasets
@@ 5.3. Discussion
However, there are limitations. The assumption that the attack type is one of the known types in Experiment 1 may not hold in all real-world scenarios. Additionally, the grid search in Experiment 2 introduces computational overhead, which could be mitigated with more efficient optimization techniques.
Future research could focus on dynamically adapting to a wider variety of attacks, expanding the framework to include additional medical imaging datasets, and exploring the integration of uncertainty-reducing adversarial training techniques. Additionally, developing novel scoring functions [52], weighting strategies [53], and entropy reweighting methods [54] could further enhance the efficiency of CP.

### v2

@@ 5.2.2. Results for RQ3
In the third experiment, an unknown and potentially adversarial attack is applied to the test data. The prediction set size corresponds to the attacker’s payoff and the defender’s cost, framing the interaction as a zero-sum game. To develop an optimal solution for the defender in a strategic manner, we preserve an additional evaluation set, which is exchangeable with both calibration and test sets. We use the prediction set size evaluated on this evaluation set as the estimated payoff matrix and derive the Nash equilibrium to determine the defender’s optimal strategy. This strategy is then applied during the test phase. Due to the unstable coverage exhibited by the VRCP-based CP methods on certain defensive models or datasets in RQ1 and RQ2, we restrict our experiments in RQ3 to only the APS and RSCP methods.
@@
The payoff matrix in this scenario has a shape of $6\times 5$, accounting for the five defensive strategies (including $f_{\text{max}}$ and $f_{\text{min}}$) and the five attack types. Furthermore, the payoff matrix describes the size of the prediction set corresponding to different defense methods under various attack methods. We divided the prediction set into two parts, one for evaluation and the other for testing. The evaluation payoff matrix (where the size of the prediction set is considered as the payoff object, with the attacker aiming to maximize it and the defender aiming to minimize it) is designed to calculate the optimal attack strategy and optimal defense strategy at the Nash equilibrium. To verify the effectiveness of our method, we computed the payoff matrix on both the evaluation set and the test set. Additionally, we incorporated a uniform strategy into the test set, where each defense model randomly selects and computes the size of the prediction set in equal proportions. All box plots illustrate the size of the prediction set for various defense models under the optimal attack strategy found on the evaluation set.
Figure 1. RQ3: Estimated and True Payoff Matrices for OrganAMNIST (APS)
Figure 2. RQ3: Estimated and True Payoff Matrices for OrganAMNIST (RSCP)
Figure 3. RQ3: Estimated and True Payoff Matrices for PathMNIST (APS)
Figure 4. RQ3: Estimated and True Payoff Matrices for PathMNIST (RSCP)
Figure 5. RQ3: Estimated and True Payoff Matrices for TissueMNIST (APS)
Figure 6. RQ3: Estimated and True Payoff Matrices for TissueMNIST (RSCP)
Figure 7. RQ3: Comparison of strategies across different datasets and CP methods.
From the reward matrix depicted in Figure 2-6, it can be discerned that, in comparison to the APS method, the RSCP generally yields larger prediction sets. This observation is supported by the findings of RQ1 and RQ2, which attribute this to the higher coverage achieved by RSCP through the utilization of more stable classification points (CPs). Furthermore, when applying the RSCP method on the PathMNIST dataset, the constructed Nash equilibrium scenario involves both the defender and the attacker employing multiple defense and attack models to reach a zero-sum game process. Consequently, in the box plot shown in Figure 7, the Nash defense model exhibits a smaller prediction set compared to other methods. In contrast, for other datasets, achieving the Nash equilibrium only requires a single strategy, which explains why the Nash defense model in Figure 7 appears nearly identical to one of the defense models. Additionally, except for the PathMNIST dataset where the two CP methods exhibit some inconsistency, the performance differences among various defense models are relatively minor across other datasets.
@@ 5.3. Discussion
However, there are limitations. The premise in RQ1, which assumes that attack types are limited to known types, may not hold true in all real-world scenarios. To address RQ2 and RQ3, we have simplified strategy selection in our current work by restricting both attackers and defenders to a discrete action space. Furthermore, for RQ3, we allow the use of weighted combinations of defense models to form mixed strategies, thereby eliminating the need for exhaustive grid search for optimal weights. This design alleviates computational overhead to a certain extent, making the method more feasible for large-scale datasets or more sophisticated adversarial strategies.
Future work may explore dynamic adaptation to a broader range of attacks, extend the framework to other medical imaging datasets, and investigate the integration of uncertainty-reducing adversarial training methods [56, 57, 58] to further enhance CP efficiency.